the width of a rectangle is 2m less than the length. The area is 48 m2. find the dimensions of the rectangle.
Answer by JulietG(1812)

Length * Width = Area
L * W = 48
W = L-2
Substitute the value of W from the second equation into the first.
L * (L-2) = 48
L = 8
Therefore, W = 8-2 or 6
.
L*W
8*6 = 48